The details of a share are given. The recent dividend it has distributed is D0=2 €, equity cost ks=13%. The dividend is expected to grow steadily at a rate of 6%. Calculate what will be the share price 2 years from today?
Answer #1

Do = 2 Di = 2x 1006 = 2.12 02 = 2012 X 1-06 = 2-25 D3= 2-25 X 1.06 = 2.38 Price en , Re=13%. g = 60% second year = D3 Re- g =